OpenStudy (anonymous):

If you were to solve the following system by substitution, what would be the best variable to solve for and from what equation? 3x+6y=9 2x-10y=13 A. y, in the second equation B. y, in the first equation C. x, in the first equation D. x, in the second equation

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

3x+6y=9 x+2y=3 x=3-2y what about the others?

OpenStudy (anonymous):

what do you mean? dont think i understand

OpenStudy (anonymous):

x from 1 eq to 2

OpenStudy (anonymous):

so C? @HARSH123

OpenStudy (anonymous):

d

OpenStudy (anonymous):

ohhh ok, my bad. thank you.